# Efficient Estimation of Relaxed Model Parameters for Robust UAV Trajectory Optimization

### v1

@@ Abstract
Online trajectory optimization and optimal control methods are crucial for enabling sustainable unmanned aerial vehicle (UAV) services, such as agriculture, environmental monitoring, and transportation, where available actuation and energy are limited. However, optimal controllers are highly sensitive to model mismatch, which can occur due to loaded equipment, packages to be delivered, or pre-existing variability in fundamental structural and thrust-related parameters. To circumvent this problem, optimal controllers can be paired with parameter estimators to improve their trajectory planning performance and perform adaptive control. However, UAV platforms are limited in terms of onboard processing power, oftentimes making nonlinear parameter estimation too computationally expensive to consider. To address these issues, we propose a relaxed, affine-in-parameters multirotor model along with an efficient optimal parameter estimator. We convexify the nominal Moving Horizon Parameter Estimation (MHPE) problem into a linear-quadratic form (LQ-MHPE) via an affine-in-parameter relaxation on the nonlinear dynamics, resulting in fast quadratic programs (QPs) that facilitate adaptive Model Predictve Control (MPC) in real time. We compare this approach to the equivalent nonlinear estimator in Monte Carlo simulations, demonstrating a decrease in average solve time and trajectory optimality cost by 98.2% and 23.9-56.2%, respectively.
@@ I Introduction
However, for MPC to demonstrate optimal behavior, the dynamics model that it employs must accurately reflect the system that is being controlled. Since MPC is sensitive to modeling errors, and performing system identification beforehand can be impractical, it may be necessary to implement modifications that improve its robustness to model uncertainty. Model mismatch can occur in drone delivery, where the system’s aerodynamic and inertial properties dramatically change before and after shipment. Other applications include agricultural crop spraying and environmental monitoring, where various attachments (e.g., sprayers, cameras, other sensors) introduce uncertainty or variability in model parameters. Although this issue can be addressed by pairing MPC with a parameter estimator (see, e.g., [10]), nonlinearities and limited onboard computing power may make online estimation challenging.
Modeling nonlinear systems with linear dynamics enables the usage of mature techniques from linear control, estimation, and optimization literature. The conventional strategy is to repeatedly linearize the nominal nonlinear dynamics about the system’s evolving state. Although these linear approximations are oftentimes serviceable, the linearization process introduces additional algorithmic complexity and approximation error. On the other hand, developments in analytical Koopman operators for rigid body dynamics [11, 12] have shown promise in producing persistently accurate linear representations of nonlinear systems. However, the literature of these analytical Koopman liftings are only interested in dynamics that are linear with respect to their states, and do not provide sufficient insight on obtaining dynamics that are affine in parameters. Fortunately, if the system is nonlinear in parameters solely due to multiplicative coupling, a simple change of variables can relax the model without the need for iterative re-linearization [13, 14].
@@
To demonstrate the effectiveness of our proposed estimator, we run Monte Carlo simulations across two different UAV models with significantly varying model parameters, multiple numerical optimization solvers, and various randomized initial conditions. Specifically, we compare nonlinear MHPE (NMHPE) with the convexified LQ-MHPE and show that LQ-MHPE runs with faster solve times and adaptively updates the model used in the NMPC to produce comparable trajectory optimality. We thereby show that LQ-MHPE enables adaptive NMPC with sufficient speed and accuracy to enhance UAV performance in sustainability applications.
@@ III Nonlinear Model Predictve Control
NMPC is a powerful optimal control strategy for systems such as (1) that optimizes performance metrics while explicitly handling constraints. For drone delivery, NMPC is especially effective for finding local energy-efficient trajectories. At every time step, NMPC solves a trajectory optimization problem over a finite future time horizon to find a sequence of future control inputs and applies the first control input in that sequence. This trajectory optimization can be formulated as the following NLP:
@@ V-B Affine-in-Parameter Relaxation
TABLE I: Crazyflie and Fusion 1 Model Parameters.
Parameter Description Crazyflie Fusion 1 Units
$\mu$ Total mass $22.70\mathrm{e}-$ 2 $12.50\mathrm{e}-$ 1 kg
$I_{xx}$ Mass moment of inertia $51.44\mathrm{e}-$ 5 $44.27\mathrm{e}-$ 4 kg m2
$I_{yy}$ Mass moment of inertia $51.40\mathrm{e}-$ 5 $46.09\mathrm{e}-$ 4 kg m2
$I_{zz}$ Mass moment of inertia $52.17\mathrm{e}-$ 5 $31.50\mathrm{e}-$ 3 kg m2
TABLE II: Simulation Trial Parameters.
Description Crazyflie Fusion 1
Parameter lower bound factor $0.5$ $0.5$
Parameter upper bound factor $1.5$ $1.5$
Process noise lower bound $-2.5$ $-2.5$
Process noise upper bound $2.5$ $2.5$
Initial position lower bound $-5.0$ $-10$
Initial position upper bound $5.0$ $10$
Initial linear vel. lower bound $-2.5$ $-5.0$
Initial linear vel. upper bound $2.5$ $5.0$
Initial angular vel. lower bound $-2.5$ $-5.0$
Initial angular vel. upper bound $2.5$ $5.0$
@@ VI-D Simulation Environment
Best-case solve time $42.43\mathrm{e}-$ 4 $33.34\mathrm{e}-$ 3 -- $42.33\mathrm{e}-$ 4 $32.96\mathrm{e}-$ 3 --
Average solve time $31.26\mathrm{e}-$ 3 $27.02\mathrm{e}-$ 2 -- $47.80\mathrm{e}-$ 4 $24.32\mathrm{e}-$ 2 --
Worst-case solve time $24.22\mathrm{e}-$ 2 $06.51\mathrm{e}-$ 0 -- $24.90\mathrm{e}-$ 2 $03.92\mathrm{e}-$ 0 --
@@
VI-E Results
For every test case, that is, every control scheme for every quadrotor model, we simulate 1,000 trials subject to random initial conditions and parameter values. Every trial is simulated for 10 seconds, with 0.02-second time steps. We use CasADi[29] to implement NMPC and MHPE, interfacing with Ipopt[30] for NMPC and NMHPE and OSQP[31] for LQ-MHPE. The convergence tolerances for Ipopt and OSQP are set to 1.0e-8 while their maximum number of iterations are set to their default values. All simulations are run on an AMD Ryzen 7 7840U 5.1 GHz processor.
Figure 3 shows an example solution from NMPC with and without accurate model parameters, which highlights its inability to handle significant model uncertainty. Figure 4 shows the norm of the 3D-position error for all 1,000 trials using NMPC paired with LQ-MHPE converging to a small value for the large majority of simulations, which signifies its robustness to significant model uncertainty. Table III and Figure 5 show the resulting MHPE solve times and recorded optimal trajectory costs from the simulation trials. These results show that the affine-in-parameter convexification of the MHPE problem yields significant computational improvements. For the Crazyflie trials, NMHPE had an average solve time of 0.07 seconds. On the other hand, the Crazyflie LQ-MHPE demonstrated an average runtime of around 1 millisecond, indicating a decrease by 98.2%. The Fusion 1 NMHPE and LQ-MHPE performed similarly to their Crazyflie counterparts, with NMHPE and LQ-MHPE yielding 0.0432 and 0.000780 seconds for their mean solve times, respectively. Again, the LQ-MHPE demonstrated an order-of-magnitude decrease in average compute time by 98.2%.
@@ VI-E Results
Although LQ-MHPE’s forward Euler integration tends to degrade in numerical accuracy compared to that of RK4 over long time steps, the computational speed of LQ-MHPE enables extremely fast (and therefore more accurate) sampling rates. The observed mean compute times for LQ-MHPE are orders of magnitude smaller than the simulated 0.02-second sampling time with minimal tuning of solver parameters, suggesting that these real-time solve rates are also achievable on board when the QP solver is optimized for the hardware of choice.
@@ VII Conclusion and Future Work
This work can be expanded in multiple directions. The first is to compare LQ-MHPE with simpler linear parameter estimators–such as the Kalman filter–to analyze the trade-offs between computational cost and constraint satisfaction. Second, there is room to explore the performance and tuning requirements of LQ-MHPE when paired with a state estimator, as opposed to assuming idealized state and disturbance observations. Finally, recent work on analytical Koopman operators for systems with rigid body dynamics can be combined with the affine-in-parameter relaxation to perform simultaneous linear state and parameter estimation.
The source code can be found at https://github.com/Herpderk/parameter-affine-relaxation.

### v2

@@ Abstract
Online trajectory optimization and optimal control methods are crucial for enabling sustainable unmanned aerial vehicle (UAV) services, such as agriculture, environmental monitoring, and transportation, where available actuation and energy are limited. However, optimal controllers are highly sensitive to model mismatch, which can occur due to loaded equipment, packages to be delivered, or pre-existing variability in fundamental structural and thrust-related parameters. To circumvent this problem, optimal controllers can be paired with parameter estimators to improve their trajectory planning performance and perform adaptive control. However, UAV platforms are limited in terms of onboard processing power, oftentimes making nonlinear parameter estimation too computationally expensive to consider. To address these issues, we propose a relaxed, affine-in-parameters multirotor model along with an efficient optimal parameter estimator. We convexify the nominal Moving Horizon Parameter Estimation (MHPE) problem into a linear-quadratic form (LQ-MHPE) via an affine-in-parameter relaxation on the nonlinear dynamics, resulting in fast quadratic programs (QPs) that facilitate adaptive Model Predictve Control (MPC). This makes real-time applications more feasible. We compare this approach to the equivalent nonlinear estimator in Monte Carlo simulations, demonstrating a decrease in average solve time by 98.2% and trajectory optimality cost by 23.9-56.2%.
@@ I Introduction
However, for MPC to demonstrate optimal behavior, the dynamics model that it employs must accurately reflect the system that is being controlled. Since MPC is sensitive to modeling errors, and performing system identification beforehand can be impractical, it may be necessary to implement modifications that improve its robustness to model uncertainty. Model mismatch can occur in drone delivery, where the system’s aerodynamic and inertial properties dramatically change before and after shipment. Other applications include agricultural crop spraying and environmental monitoring, where various attachments (e.g., sprayers, cameras, other sensors) introduce uncertainty or variability in model parameters. Although this issue can be addressed by pairing MPC with a parameter estimator (see, e.g., [10]), nonlinearities and limited onboard computing power make online estimation challenging.
Modeling nonlinear systems with linear dynamics enables the usage of mature techniques from linear control, estimation, and optimization literature. The conventional strategy is to repeatedly linearize the nominal nonlinear dynamics about the system’s evolving state. Although these linear approximations are oftentimes serviceable, the linearization process introduces additional algorithmic complexity and approximation error. Fortunately, if the system is nonlinear in parameters solely due to multiplicative coupling, a simple change of variables can relax the model without the need for iterative re-linearization [11, 12].
@@
To demonstrate the effectiveness of our proposed estimator, we run Monte Carlo simulations across two different UAV models with significantly varying model parameters, multiple numerical optimization solvers, and various randomized initial conditions. Specifically, we compare nonlinear MHPE (NMHPE) with the convexified LQ-MHPE and show that LQ-MHPE runs with faster solve times and adaptively updates the model used in the NMPC to produce comparable, and in some cases better, trajectory optimality. We thereby show that LQ-MHPE enables adaptive NMPC with sufficient speed and accuracy to enhance UAV performance in sustainability applications.
@@ III Nonlinear Model Predictve Control
NMPC is a powerful optimal control strategy for systems such as (1) that optimizes performance metrics while explicitly handling constraints. For drone delivery, NMPC is especially effective for finding local energy-efficient trajectories. At every time step, NMPC solves a trajectory optimization problem over a finite time horizon to find a sequence of future control inputs and applies the first input in said sequence. This trajectory optimization can be formulated as the following NLP:
@@ V-A Nonlinear Formulation
TABLE I: Crazyflie and Fusion 1 Model Parameters.
Parameter Description Crazyflie Fusion 1 Units
$\mu$ Total mass $22.70\mathrm{e}-$ 2 $12.50\mathrm{e}-$ 1 kg
$I_{xx}$ Mass moment of inertia $51.44\mathrm{e}-$ 5 $44.27\mathrm{e}-$ 4 kg m2
$I_{yy}$ Mass moment of inertia $51.40\mathrm{e}-$ 5 $46.09\mathrm{e}-$ 4 kg m2
$I_{zz}$ Mass moment of inertia $52.17\mathrm{e}-$ 5 $31.50\mathrm{e}-$ 3 kg m2
@@ VI-D Simulation Environment
VI-E Results
For every test case, that is, every control scheme for every quadrotor model, we simulate 1,000 trials subject to random initial conditions and parameter values. Every trial is simulated for 10 seconds, with 0.02-second time steps. We use CasADi[27] to implement NMPC and MHPE, interfacing with Ipopt[28] for NMPC and NMHPE and OSQP[29] for LQ-MHPE. The convergence tolerances for Ipopt and OSQP are set to 1.0e-8 while their maximum number of iterations are set to their default values. Although our simulations are run on an AMD Ryzen 7 7840U 5.1 GHz processor that may not be feasible for onboard computing, the relative differences in performance should still carry significance on resource-constrained systems. Additionally, we use NMPC as the baseline because it is the most optimistic, albeit computationally expensive, controller. In practice, a suboptimal linear MPC could be run onboard as a QP to achieve real-time computation.
TABLE II: Simulation Trial Parameters.
Description Crazyflie Fusion 1
Parameter lower bound factor $0.5$ $0.5$
Parameter upper bound factor $1.5$ $1.5$
Process noise lower bound $-2.5$ $-2.5$
Process noise upper bound $2.5$ $2.5$
Initial position lower bound (m) $-5.0$ $-10$
Initial position upper bound (m) $5.0$ $10$
Initial linear vel. lower bound (m/s) $-2.5$ $-5.0$
Initial linear vel. upper bound (m/s) $2.5$ $5.0$
Initial angular vel. lower bound (rad/s) $-2.5$ $-5.0$
Initial angular vel. upper bound (rad/s) $2.5$ $5.0$
Sampling time (s) $0.02$ $0.02$
@@ VI-E Results
Best-case solve time (s) $42.43\mathrm{e}-$ 4 $33.34\mathrm{e}-$ 3 -- $42.33\mathrm{e}-$ 4 $32.96\mathrm{e}-$ 3 --
Average solve time (s) $31.26\mathrm{e}-$ 3 $27.02\mathrm{e}-$ 2 -- $47.80\mathrm{e}-$ 4 $24.32\mathrm{e}-$ 2 --
Worst-case solve time (s) $24.22\mathrm{e}-$ 2 $06.51\mathrm{e}-$ 0 -- $24.90\mathrm{e}-$ 2 $03.92\mathrm{e}-$ 0 --
@@
Figure 3 shows example NMPC simulation trials with and without accurate model parameters, which highlights its inability to handle significant model uncertainty. Figure 4 shows the norm of the 3D-position error for all 1,000 trials using NMPC paired with LQ-MHPE converging to a small value for the large majority of simulations, which signifies its robustness to significant model uncertainty. Table III and Figure 5 show the resulting MHPE solve times and recorded optimal trajectory costs from the simulation trials. These results show that the affine-in-parameter convexification of the MHPE problem yields significant computational improvements. For the Crazyflie trials, NMHPE had an average solve time of 0.07 seconds. On the other hand, the Crazyflie LQ-MHPE demonstrated an average runtime of around 1 millisecond, indicating a decrease by 98.2%. The Fusion 1 NMHPE and LQ-MHPE performed similarly to their Crazyflie counterparts, with NMHPE and LQ-MHPE yielding 0.0432 and 0.000780 seconds for their mean solve times, respectively. Again, the LQ-MHPE demonstrated an order-of-magnitude decrease in average compute time by 98.2%.
@@
Although LQ-MHPE’s forward Euler integration tends to degrade in numerical accuracy compared to that of RK4 over long time steps, the computational speed of LQ-MHPE enables extremely fast (and therefore more accurate) sampling rates. The observed mean compute times for LQ-MHPE are orders of magnitude smaller than the simulated 0.02-second sampling time with minimal tuning of solver parameters, suggesting that these real-time solve rates are also achievable on board when the QP solver is optimized for the processor of choice.
@@ VII Conclusion and Future Work
This work can be expanded in multiple directions. The first is onboard hardware implementation–it has been demonstrated that linear MPC can be run in real time on microcontrollers [20], and the proposed numerical methods should be naturally extendable to LQ-MHPE. The second is to compare LQ-MHPE with simpler linear parameter estimators–such as the Kalman filter–to analyze the trade-offs between computational cost and constraint satisfaction. Finally, there is room to investigate the performance and tuning of the control/parameter estimation scheme in the presence of state estimation (as opposed to idealized state measurements).
The source code can be found at https://github.com/EASEL-UCI/parameter-affine-relaxation.